Find the P value for the indicated hypothesis test with the given standardized test statistic z Decide whether to reject Ho for the given level of significance ? Right-tailed test with test statistic z-1.55 and ?-0.04 P-value(Round to four decimal places as needed.) State your conclusion O Reject Ho O Fail to reject Ho
Answer #1

Solution :

Given that,

z = 1.55 and \alpha = 0.04

This is the right tailed test .

P(Z > 1.55) = 1 - P(Z < 1.55) = 1 - 0.9394 = 0.0606

P-value = 0.0606

P-value > \alpha

Fail t0 reject H0